# Lattice-induced sound trapping in biperiodic metasurfaces of acoustic resonators

### v3

@@ Appendix C Acoustic T-matrix of a sphere
For an isolated homogeneous sphere of diameter $a$, the entries of the acoustic T-matrix read as $T_{\ell,m,\ell^{\prime},m^{\prime}}=\mathfrak{a}_{\ell}\delta_{\ell,\ell^{\prime}}\delta_{m,m^{\prime}}$ where $\delta_{\ell,\ell^{\prime}}$ is the Kroneker symbol, and the coefficients $\mathfrak{a}_{\ell}$ are given as [83]
@@
We can rewrite the inverse coefficient in the absence of material losses (absorption), i.e., when the material parameters are real-valued, as
@@
Here, we have used the definition of the spherical Hankel function $h^{(1)}_{\ell}(\xi)=j_{\ell}(\xi)+\mathrm{i}y_{\ell}(\xi)$ where $y_{\ell}(\xi)$ is the spherical Bessel function of the second kind. Hence,
@@
$\Psi^{(3)}_{\ell,m}(\omega;-\mathbf{R})=\sum_{\ell^{\prime},m^{\prime}}\mathcal{A}^{(3)}_{\ell^{\prime},m^{\prime},\ell,m}(\omega;-\mathbf{R})\Psi^{(1)}_{\ell^{\prime},m^{\prime}}(\omega;\bm{0})\,.$ (38)
@@
$\begin{aligned} \widetilde{b}_{\ell,m}=&b_{\ell,m}+\\ &\sum_{\ell^{\prime},m^{\prime}}\sum_{{\bf R}\neq\bm{0}}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})a_{{\bf R},\ell^{\prime},m^{\prime}}^{\rm eff}\,.\end{aligned}$ (40)
Substituting Eq. (8) into Eq. (40) gives us the following
@@
$\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\mathbf{k}_{\parallel})=\sum_{{\bf R}\neq\bm{0}}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{R}}\,.$ (42)
@@
$\begin{aligned} a_{\ell,m}^{\rm eff}=&\sum_{\ell^{\prime},m^{\prime}}T_{\ell,m,\ell,m^{\prime}}(\omega)\\ &\times\left[b_{\ell^{\prime},m^{\prime}}+\sum_{\ell^{\prime\prime},m^{\prime\prime}}\Sigma^{\prime}_{\ell^{\prime},m^{\prime},\ell^{\prime\prime},m^{\prime\prime}}(\omega,\mathbf{k}_{\parallel})a_{\ell^{\prime\prime},m^{\prime\prime}}^{\rm eff}\right]\,.\end{aligned}$ (43)
@@ Appendix E Derivation of the selection rules for $\mathbf{k}_{\parallel}=\bf{0}$
Let us derive selection rules (12) for lattice-induced coupling in the case of normal incidence $\mathbf{k}_{\parallel}=\bf{0}$. For this purpose, we consider off-diagonal entries of the inverse effective T-matrix that are $\left(-\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}\right)$ if the T-matrix of an isolated resonator is diagonal. The definition of $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}$ for $\mathbf{k}_{\parallel}=\bm{0}$ is [see (42)]
$\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\sum_{{\bf R}\neq\bm{0}}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\,.$ (44)
First, we can note that for a lattice in the $xy$ plane, each term of the lattice sum is proportional to $P_{\ell+\ell^{\prime}-2s}^{m-m^{\prime}}(0)$ where $s$ is an integer. $P_{\ell+\ell^{\prime}-2s}^{m-m^{\prime}}(0)$ is zero unless $(\ell+\ell^{\prime}-m+m^{\prime})$ is an even number [64]. However, we can formulate more specific selection rules by analyzing the behavior of the lattice sum under the symmetry operations.
E.1: Inversion symmetry
First, we consider a lattice that has symmetry with respect to the inversion $\mathbf{R}\to-\mathbf{R}$. After the inversion, (44) becomes the following
$\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\sum_{{\bf R}\neq\bm{0}}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{R})\,.$ (45)
The translation coefficients of scalar spherical waves obey the following property [84]
$\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{R})=(-1)^{\ell+\ell^{\prime}}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\,.$ (46)
However, since the lattice remains the same upon inversion, the lattice sums must be equal $\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}=\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}$, which implies that $(\ell+\ell^{\prime})$ is an even number. This can be true only if $\ell$ and $\ell^{\prime}$ are equal modulo 2, i.e., they have the same parity. Note that the parity of degree $\ell$ determines the parity of a spherical multipole under the inversion of coordinates.
@@
Similarly, one can consider the rotations around the $z$ axis by angle $\varphi=2\pi/\mathfrak{n}$. According to (32), the lattice sums upon the rotation can be written as
$\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\mathrm{e}^{\mathrm{i}2\pi(m-m^{\prime})/\mathfrak{n}}\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\bm{0})\,.$ (47)
One can clearly see that the lattice sum is conserved if $(m-m^{\prime})=s\mathfrak{n}$ where $s$ is an integer; that is, $m$ and $m^{\prime}$ are equal modulo $\mathfrak{n}$.
@@ Appendix F Calculation of the lattice field sums
Let us calculate lattice sums (14) in the far-field approximation $kr\gg 1$. First, we will use the integral expansion of the singular scalar spherical wave into plane waves [85]
@@
where $A$ is the area of a unit cell, $\mathbf{G}$ is a vector of the reciprocal lattice and $\delta^{(2)}$ is the 2D Dirac delta function, which eliminates the integration. After combining Eqs. (48), (49), and (14), we can express the lattice sum as
@@
In the following, we calculate the real parts of the lattice sums that are present in (25). Combining Eqs. (44), (32), and (15), one can obtain that
$\begin{aligned} &\Sigma_{\ell,0,\ell,0}^{\prime}(\omega,\mathbf{0})=\sqrt{4\pi}(2\ell+1)\sum\limits_{\begin{subarray}{c}q=0\\ q\ \text{even}\end{subarray}}^{2\ell}\mathrm{i}^{q}\sqrt{2q+1}\times\\ &\times\lim\limits_{\mathbf{r}\to\mathbf{0}}\left[\Sigma_{q,0}(\omega,\mathbf{0};\mathbf{r})-\Psi^{(3)}_{q,0}(\omega;\mathbf{r})\right]\begin{pmatrix}\ell&\ell&q\\ 0&0&0\end{pmatrix}^{2}\,.\end{aligned}$ (51)
@@
where $P_{\ell}^{0}(1)\equiv 1$. For the second term in (51), we use (28) and obtain that
@@
while $\lim\limits_{\mathbf{r}\to\mathbf{0}}\operatorname{Im}\left[\Psi^{(3)}_{q,0}(\omega;\mathbf{r})\right]$ does not exist. Inserting Eqs. (52) and (53) into Eq. (51) and using the properties of the Wigner’s 3-j symbols [81], we obtain that
@@
$\begin{aligned} \operatorname{Re}\left[\Sigma_{\ell,0,\ell^{\prime},0}^{\prime}(\omega,\mathbf{0})\right]=\sqrt{(2\ell+1)(2\ell^{\prime}+1)}\mathrm{i}^{\ell^{\prime}-\ell}\frac{\lambda_{\rm b}^{2}}{2\pi A}\cdot\\ \cdot\underbrace{\sum\limits_{\begin{subarray}{c}q=\ell^{\prime}-\ell\\ q\ \text{even}\end{subarray}}^{\ell+\ell^{\prime}}(2q+1)\begin{pmatrix}\ell&\ell^{\prime}&q\\ 0&0&0\end{pmatrix}^{2}}_{=1}\,.\end{aligned}$ (55)
@@ Appendix H Relationship between the BIC conditions
In this section, we show the equivalence of BIC conditions (21) and (26) for matrix (25). First, we introduce the monopole and quadrupole T-matrix elements of a sphere modified due to the monopole-monopole and quadrupole-quadrupole coupling, respectively
@@
Considering Eqs. (36) and (56), one can see that for $L<\lambda_{\rm b}$, the real parts of (57) are related to each other as
@@
The determinant of (25) being zero implies that
@@
Due to (56), we also have for $L<\lambda_{\rm b}$ that
@@
Hence, the solution to (59) reads as
@@
Indeed, considering (58) and (60), we can write that $1/T_{\rm MM}^{\rm mod}=A+\mathrm{i}B$, $1/T_{\rm QQ}^{\rm mod}=5A+\mathrm{i}C$, and $\Sigma^{\prime}_{\rm MQ}=\sqrt{5}A+\mathrm{i}D$, where $A$, $B$, $C$, and $D$ are real-valued numbers, and plug them into Eq. (59). Equating the real and imaginary parts of the left and right sides of the equation yields the following system of equations
@@
The solution is $B=D/\sqrt{5}$ and $C=D\sqrt{5}$, and we obtain (61). If an eigenmode condition (61) is satisfied for $\omega\in\mathbb{R}$, then the lattice mode is a BIC.
If we rewrite Eq. (20) using (25) and (57), we obtain that
@@
Under BIC condition (61), we immediately obtain (26), i.e., $\left(a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}\right)=0$ corresponding to the far-field intensity $\mathcal{P}_{\pm}=0$. Note that, for a normally incident plane wave of amplitude $p_{0}$, Eq. (10) under the BIC condition gives that $\left(a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}\right)\propto p_{0}$ and consequently $\mathcal{P}_{\pm}\neq 0$. Hence, the BIC cannot be excited by a normally incident plane wave.
@@ Appendix I Odd $\Gamma$-point BIC in the dipole-octupole approximation
In the main text, we consider the BIC formed by the zonal multipoles of even parity. Similarly, we can consider the BIC formed by the multipoles with odd parity, the dipole and octupole. The frequency and lattice constant of the odd-parity BIC can also be determined by Eq. (22) with the following inverse effective T-matrix
@@
where, for a sphere, $T_{\rm DD}\equiv\mathfrak{a}_{1}$ and $T_{\rm OO}\equiv\mathfrak{a}_{3}$ [see Eq. (34)]. Here, the lattice sums are $\Sigma^{\prime}_{\rm DD}\equiv\Sigma^{\prime}_{\rm 1,0,1,0}(\omega,\bm{0}),\Sigma^{\prime}_{\rm DO}\equiv\Sigma^{\prime}_{\rm 0,0,3,0}(\omega,\bm{0}),$ and $\Sigma^{\prime}_{\rm OO}\equiv\Sigma^{\prime}_{\rm 3,0,3,0}(\omega,\bm{0})$ [see Eq. (42)]. Using the same approach as in Appendix H, from the condition $\left[\det\mathbf{T}^{-1}_{\rm eff}(\omega)=0\right]$ with $\omega\in\mathbb{R}$, we derive the constraint on the multipole moments of the odd BIC
$\sqrt{3}a^{\rm eff}_{\rm D}-\sqrt{7}a^{\rm eff}_{\rm O}=0\,.$ (65)
This condition provides $\mathcal{P}_{+}=\mathcal{P}_{-}=0$ that corresponds to the BIC regime. Figure 7 presents the same content as Fig. 3 but for the odd BIC.
@@
$\begin{aligned} \widetilde{b}_{i,\ell,m}(\omega)=&b_{i,\ell,m}(\omega)\\ &+\sum_{\ell^{\prime},m^{\prime}}\sum_{j\neq i}\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{r}_{i}-\mathbf{r}_{j})a_{j,\ell^{\prime},m^{\prime}}(\omega)\end{aligned}$ (68)
and substitute them into Eq. (7) for each scatterer to obtain the following system of linear equations
$\mathbf{a}_{i}=\mathbf{T}_{i}\left[\mathbf{b}_{i}+\sum_{j\neq i}\bm{\mathcal{A}}^{(3)}_{ij}\mathbf{a}_{j}\right]\,$, (69)
where $\mathbf{a}_{i}\equiv\mathbf{a}_{i}(\omega)$, $\mathbf{b}_{i}\equiv\mathbf{b}_{i}(\omega)$, $\mathbf{T}_{i}\equiv\mathbf{T}_{i}(\omega)$ is the acoustic T-matrix of the $i$-th scatterer and $\bm{\mathcal{A}}^{(3)}_{ij}$ is the matrix that contains the translation coefficients $\mathcal{A}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{r}_{i}-\mathbf{r}_{j})$.

### v4

@@ Appendix C Acoustic T-matrix of a sphere
For an isolated homogeneous sphere of diameter $a$, the entries of the acoustic T-matrix read as $T_{\ell,m,\ell^{\prime},m^{\prime}}=\mathfrak{a}_{\ell}\delta_{\ell,\ell^{\prime}}\delta_{m,m^{\prime}}$, where $\delta_{\ell,\ell^{\prime}}$ is the Kronecker symbol, and the coefficients $\mathfrak{a}_{\ell}$ are given as [88]
@@
We can rewrite the inverse coefficient in the absence of material losses (absorption), i.e., when the material parameters are real valued, as
@@
Here, we have used the definition of the spherical Hankel function $h^{(1)}_{\ell}(\xi)=j_{\ell}(\xi)+\mathrm{i}y_{\ell}(\xi)$, where $y_{\ell}(\xi)$ is the spherical Bessel function of the second kind. Hence,
@@
$\Psi^{(3)}_{\ell,m}(\omega;-\mathbf{R})=\sum_{\ell^{\prime},m^{\prime}}\mathcal{C}^{(3)}_{\ell^{\prime},m^{\prime},\ell,m}(\omega;-\mathbf{R})\Psi^{(1)}_{\ell^{\prime},m^{\prime}}(\omega;\bm{0})\,.$ (39)
@@
$\begin{aligned} \widetilde{b}_{\ell,m}=&b_{\ell,m}\\ &+\sum_{\ell^{\prime},m^{\prime}}\sum_{{\bf R}\neq\bm{0}}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})a_{{\bf R},\ell^{\prime},m^{\prime}}^{\rm eff}\,.\end{aligned}$ (41)
Substituting Eq. (8) into Eq. (41) gives us the following:
@@
$\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\mathbf{k}_{\parallel})=\sum_{{\bf R}\neq\bm{0}}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{R}}\,.$ (43)
@@
$\begin{aligned} a_{\ell,m}^{\rm eff}=&\sum_{\ell^{\prime},m^{\prime}}T_{\ell,m,\ell^{\prime},m^{\prime}}(\omega)\\ &\times\left[b_{\ell^{\prime},m^{\prime}}+\sum_{\ell^{\prime\prime},m^{\prime\prime}}\Sigma^{\prime}_{\ell^{\prime},m^{\prime},\ell^{\prime\prime},m^{\prime\prime}}(\omega,\mathbf{k}_{\parallel})a_{\ell^{\prime\prime},m^{\prime\prime}}^{\rm eff}\right]\,.\end{aligned}$ (44)
@@ Appendix E Derivation of the selection rules for $\mathbf{k}_{\parallel}=\bf{0}$
Let us derive selection rules (12), i.e., determine nonzero $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}$ for $\mathbf{k}_{\parallel}=\bf{0}$ given by [see Eq. (43)]
$\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\sum_{{\bf R}\neq\bm{0}}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\,.$ (45)
1. Inversion symmetry
First, we consider a lattice that has symmetry with respect to the inversion $\mathbf{R}\to-\mathbf{R}$. After the inversion, Eq. (45) becomes the following:
$\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\sum_{{\bf R}\neq\bm{0}}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{R})\,.$ (46)
Since the parity of degree $\ell$ determines the parity of a spherical multipole under the inversion $\Psi^{(n)}_{\ell,m}(\omega;-\mathbf{r})=(-1)^{\ell}\Psi^{(n)}_{\ell,m}(\omega;\mathbf{r})$, the translation coefficients of scalar spherical waves obey the following property [89]:
$\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{R})=(-1)^{\ell+\ell^{\prime}}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;-\mathbf{R})\,.$ (47)
However, since the lattice remains the same upon inversion, the lattice sums must be equal $\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}=\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}$, which implies that $(\ell+\ell^{\prime})$ is an even number. This can be true only if $\ell$ and $\ell^{\prime}$ are equal modulo 2; i.e., they have the same parity.
@@ 1. Inversion symmetry
Similarly, one can consider rotations around the $z$ axis by angle $\varphi=2\pi/\mathfrak{n}_{\rm L}$. According to Eq. (33), the lattice sums can be written upon the rotation as
$\widehat{\Sigma}^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\bm{0})=\mathrm{e}^{\mathrm{i}2\pi(m-m^{\prime})/\mathfrak{n}_{\rm L}}\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\bm{0})\,.$ (48)
One can clearly see that the lattice sum is preserved only if $(m-m^{\prime})=s\mathfrak{n}$, where $s$ is an integer; that is, $m$ and $m^{\prime}$ are equal modulo $\mathfrak{n}_{\rm L}$.
@@ Appendix F Calculation of the lattice field sums
Let us calculate lattice sums (14) in the far-field approximation $kr\gg 1$. First, we will use the integral expansion of the singular scalar spherical wave into plane waves [90],
@@
where $A$ is the area of a unit cell, $\mathbf{G}$ is a vector of the reciprocal lattice, and $\delta^{(2)}$ is the 2D Dirac delta function, which eliminates the integration. After combining Eqs. (49), (50), and (14), we can express the lattice sum as
@@
In the following, we calculate the real parts of the lattice sums that are present in Eq. (25). Combining Eqs. (45), (33), and (15), one can obtain that
$\begin{aligned} &\Sigma_{\ell,0,\ell,0}^{\prime}(\omega,\mathbf{0})=\sqrt{4\pi}(2\ell+1)\sum\limits_{\begin{subarray}{c}q=0\\ q\ \text{even}\end{subarray}}^{2\ell}\mathrm{i}^{q}\sqrt{2q+1}\\ &\times\lim\limits_{\mathbf{r}\to\mathbf{0}}\left[\Sigma_{q,0}(\omega,\mathbf{0};\mathbf{r})-\Psi^{(3)}_{q,0}(\omega;\mathbf{r})\right]\begin{pmatrix}\ell&\ell&q\\ 0&0&0\end{pmatrix}^{2}\,.\end{aligned}$ (52)
@@
where $P_{\ell}^{0}(1)\equiv 1$. For the second term in Eq. (52), we use Eq. (29) and obtain that
@@
while $\lim\limits_{\mathbf{r}\to\mathbf{0}}\operatorname{Im}\left[\Psi^{(3)}_{q,0}(\omega;\mathbf{r})\right]$ does not exist. Inserting Eqs. (53) and (54) into Eq. (52) and using the properties of the Wigner’s $3j$ symbols [86], we obtain that
@@
$\begin{aligned} \operatorname{Re}\left[\Sigma_{\ell,0,\ell^{\prime},0}^{\prime}(\omega,\mathbf{0})\right]=\sqrt{(2\ell+1)(2\ell^{\prime}+1)}\mathrm{i}^{\ell^{\prime}-\ell}\frac{\lambda_{\rm b}^{2}}{2\pi A}\\ \times\underbrace{\sum\limits_{\begin{subarray}{c}q=\ell^{\prime}-\ell\\ q\ \text{even}\end{subarray}}^{\ell+\ell^{\prime}}(2q+1)\begin{pmatrix}\ell&\ell^{\prime}&q\\ 0&0&0\end{pmatrix}^{2}}_{=1}\,.\end{aligned}$ (56)
@@ Appendix H Relationship between the BIC conditions
In this section, we show the equivalence of BIC conditions (21) and (26) for matrix (25). First, we introduce the monopole and quadrupole T-matrix elements of a sphere modified due to the monopole-monopole and quadrupole-quadrupole coupling, respectively,
@@
Considering Eqs. (37) and (57), one can see that for $L<\lambda_{\rm b}$, the real parts of Eq. (58) are related to each other as
@@
The determinant of Eq. (25) being zero implies that
@@
Due to Eq. (57), we also have for $L<\lambda_{\rm b}$ that
@@
Hence, the solution to Eq. (60) reads as
@@
Indeed, considering (59) and (61), we can write that $1/T_{\rm MM}^{\rm mod}=A+\mathrm{i}B$, $1/T_{\rm QQ}^{\rm mod}=5A+\mathrm{i}C$, and $\Sigma^{\prime}_{\rm MQ}=\sqrt{5}A+\mathrm{i}D$, where $A$, $B$, $C$, and $D$ are real-valued numbers, and plug them into Eq. (60). Equating the real and imaginary parts of the left and right sides of the equation yields the following system of equations:
@@
The solution is $B=D/\sqrt{5}$ and $C=D\sqrt{5}$, and we obtain Eq. (62). If an eigenmode condition (62) is satisfied for $\omega\in\mathbb{R}$, then the lattice mode is a BIC.
If we rewrite Eqs. (20) using (25) and (58), we obtain that
@@
Under BIC condition (62), we immediately obtain Eq. (26), i.e., $\left(a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}\right)=0$ corresponding to the far-field intensity $\mathcal{P}_{\pm}=0$. Note that, for a normally incident plane wave of amplitude $p_{0}$, Eq. (10) under the BIC condition gives that $\left(a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}\right)\propto p_{0}$ and consequently $\mathcal{P}_{\pm}\neq 0$. Hence, the BIC cannot be excited by a normally incident plane wave.
@@ Appendix I Odd $\Gamma$-point BIC in the dipole-octupole approximation
In the main text, we consider the BIC formed by the zonal multipoles of even parity. Similarly, we can consider the BIC formed by the multipoles with odd parity, the dipole and octupole. The frequency and lattice constant of the odd-parity BIC can also be determined by Eq. (22) with the following inverse effective T-matrix:
@@
where, for a sphere, $T_{\rm DD}\equiv\mathfrak{a}_{1}$ and $T_{\rm OO}\equiv\mathfrak{a}_{3}$ [see Eq. (35)]. Here, the lattice sums are $\Sigma^{\prime}_{\rm DD}\equiv\Sigma^{\prime}_{\rm 1,0,1,0}(\omega,\bm{0}),\Sigma^{\prime}_{\rm DO}\equiv\Sigma^{\prime}_{\rm 0,0,3,0}(\omega,\bm{0}),$ and $\Sigma^{\prime}_{\rm OO}\equiv\Sigma^{\prime}_{\rm 3,0,3,0}(\omega,\bm{0})$ [see Eq. (43)]. Note that additional off-diagonal terms due to $T_{\rm DO}=T_{\rm OD}\neq 0$ appear in Eq. (65) for nonspherical particles. Using the same approach as in Appendix H, we can derive Eq. (27) from condition $\left[\det\mathbf{T}^{-1}_{\rm eff}(\omega)=0\right]$ with $\omega\in\mathbb{R}$. Figure 7 presents the same content as Fig. 3, but for the odd BIC.
@@
$\begin{aligned} \widetilde{b}_{i,\ell,m}(\omega)=&b_{i,\ell,m}(\omega)\\ &+\sum_{\ell^{\prime},m^{\prime}}\sum_{j\neq i}\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{r}_{i}-\mathbf{r}_{j})a_{j,\ell^{\prime},m^{\prime}}(\omega)\end{aligned}$ (68)
and substitute them into Eq. (7) for each scatterer to obtain the following system of linear equations:
$\mathbf{a}_{i}=\mathbf{T}_{i}\left[\mathbf{b}_{i}+\sum_{j\neq i}\bm{\mathcal{C}}^{(3)}_{ij}\mathbf{a}_{j}\right]\,$, (69)
where $\mathbf{a}_{i}\equiv\mathbf{a}_{i}(\omega)$, $\mathbf{b}_{i}\equiv\mathbf{b}_{i}(\omega)$, $\mathbf{T}_{i}\equiv\mathbf{T}_{i}(\omega)$ is the acoustic T-matrix of the $i$th scatterer, and $\bm{\mathcal{C}}^{(3)}_{ij}$ is the matrix that contains the translation coefficients $\mathcal{C}^{(3)}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega;\mathbf{r}_{i}-\mathbf{r}_{j})$.
Appendix K Description of the FEM model
[…]